OpenStudy (anonymous):

To rent a certain meeting room, a college charges a reservation fee of 38.00 and an additional fee of 9.80 per hour. The chemistry club wants to spend less than 126.20 on renting the meeting room. What are the possible amounts of time for which they could rent the meeting room? Use t for the number of hours the meeting room is rented, and solve your inequality for t

OpenStudy (anonymous):

the room could be rented for 9 hours \[126.20-38.00=88.20 \] \[88.20\div9.80=9\]

OpenStudy (anonymous):

would you mind helping me with somethimg else

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

I have flvs and I need it to graduate I need help with the test n quizzes

OpenStudy (anonymous):

flv??? not sure what you mean im only a freshman in high school

OpenStudy (anonymous):

are you good in math

OpenStudy (anonymous):

b+ A- work yes

OpenStudy (nurali):

38+9.80t<=126.20 9.80t<=88.2 t<=88.2/9.80 t<=9 hours